What is the simple interest on a Php500 loan at 2.5% interest to be paid at the end of 6 months?

1
Expert's answer
2021-11-17T17:23:09-0500

"S. I=\frac {PRT} {100}"


P=500

R=2.5%

T="\frac {6}{12}=0.5"


"S. I=\frac {500×2.5×0.5}{100}"


"S. I=\frac {625}{100}"


"SI=Php6.25"